OpenStudy (anonymous):

Find a solution with theta in radians (of possible) 1. tan(theta)=947 3. 3cos(theta)=0.714 5. tan(5(theta)+7=-0.241

OpenStudy (anonymous):

please explain how you solve for these. I don't get them.

OpenStudy (anonymous):

1. theta = 1.5697 radians -this is the angle whose tan is 947 - use calculator use the tan-1 key

OpenStudy (anonymous):

1. tan(theta)947 theta = arctan(947) theta = 1.56974 3. 3cos(theta)=0.714 3cos(theta)/3=0.714/3 theta = arccos(0.238) theta = 1.33049 5. tan(5(theta))+7 = -0.241 tan(5(theta)) = -7.241 5(theta) = arctan(-7.241) theta = -1.433562/5 theta = -0.286712

OpenStudy (anonymous):

3. cos theta = 0.714/3 - use calculator cos-1 gives 1.33 radians

OpenStudy (anonymous):

5 - theta = -0.287 radians